OpenStudy (sillybutt100):

How do I find the exact value of this expression? (5)sqrt13+sqrt6 -7*sqrt13 The equation will also be in the comments

OpenStudy (sillybutt100):

\[5\sqrt{13}+\sqrt{6}-7\sqrt{13}\]

OpenStudy (anonymous):

i am not sure what you are asking, You can simplify the expression by combining the sqrt(13) so the equation becomes \[\sqrt{6}-2\sqrt{13}\] from there, can you use a calculator?

OpenStudy (sillybutt100):

I don't know what it is asking either and I don't have a book I even googled it and couldn't find anything

OpenStudy (anonymous):

i believe the simplified answer i gave before is the one you are looking for. A calculator gives you an approximate solution not an exact answer

OpenStudy (anonymous):

for example the exact solution of \[\sqrt{8}=2\sqrt{2}\]

OpenStudy (anonymous):

this is just a fancy way of asking simplify the expression as much as you can

OpenStudy (sillybutt100):

Thanks